# Distributed games with jumps: an $\alpha$-potential game Approach

### v1

@@ Abstract.
Motivated by game-theoretic models of crowd motion dynamics, this paper analyzes a broad class of distributed games with jump diffusions within the recently developed $\alpha$-potential games framework. We demonstrate that analyzing the $\alpha$-Nash equilibria is equivalent to solving finite-dimensional control problems. Beyond the viscosity and verification characterizations for the general games, we explicitly and in detail examine how spatial population distributions and interaction rules influence the structure of $\alpha$-Nash equilibria in these distributed settings, and in particular for crowd motion games.
Our theoretical results are supported by numerical implementations using policy gradient-based algorithms, further demonstrating the computational advantages of the $\alpha$-potential game framework in computing Nash equilibria for general dynamic games.
@@
Potential game, Nash equilibrium, distributed game, Hamilton–Jacobi–Bellman equation, policy-gradient algorithm, game-theoretic motion planning
@@ 1. Introduction
${\mathrm{d}}X_{i,t}$ $=b_{i}(t)u_{i,t}{\mathrm{d}}t+\sigma_{i}(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\gamma_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad t\in(0,T];\quad X_{i,0}=x_{i}\in{\mathbb{R}}^{d},$ (1.1)
@@
$J_{i}({\textbf{{u}}})\coloneqq{\mathbb{E}}\left[\int_{0}^{T}\left(\ell_{i}(u_{i,t})+\frac{1}{N-1}\sum_{j=1,j\neq i}^{N}q_{ij}K(X^{u_{i}}_{i,t}-X^{u_{j}}_{j,t})\right)\,{\mathrm{d}}t+g_{i}(X^{u_{i}}_{i,T})\right],$ (1.2)
@@
The above crowd motion game is a special class of stochastic differential games which we name as distributed games; See Section 2 for the detailed formulation of these games. The term “distributed” refers to the characteristics of the game where each player’s dynamics evolve according to a controlled stochastic process that depends only on her own control, while her objective function may depend on the joint state and control profiles of all players (see also Remark 2.1). Such a framework has been used in a variety of applications where agents interact through their objectives but evolve independently in state, including distributed control of multi-agent systems and trajectory planning [23, 42, 43, 1], transportation and routing [5, 9, 25], as well as in energy markets and smart grids [29, 41, 36, 34, 45, 32].
@@
Meanwhile, the recently introduced $\alpha$-potential game framework has shown significant promise for analyzing and solving general dynamic games, both from theoretical and algorithmic perspectives [15, 18, 17, 16, 22, 30]. For example, [15] and [17] demonstrate that computing a Nash equilibrium can be reformulated as an optimization problem involving a single $\alpha$-potential function and the analysis of the parameter $\alpha$. Furthermore, optimizing this $\alpha$-potential function has been shown in [17] to be equivalent to solving a conditional McKean–Vlasov control problem. Consequently, an $\alpha$-Nash equilibrium of the stochastic game can be characterized by an infinite-dimensional Hamilton–Jacobi–Bellman (HJB) equation. Unlike the conventional mean field game approach, which relies on weak interactions among players or takes the limit as the number of players $N\to\infty$, the $\alpha$-potential game framework can be directly applied to finite-player settings.
@@
In this paper, we adopt the $\alpha$-potential game framework to analyze distributed games with controlled jump diffusions under an open-loop setting. Within this framework, we show that the task of finding an $\alpha$-Nash equilibrium can be further reduced to solving a finite-dimensional control problem (Theorem 3.1). This reduction enables the application of standard tools, such as the viscosity solution method and the verification theorem, to characterize the $\alpha$-Nash equilibria via a finite-dimensional HJB equation (Theorems 4.1 and 4.2).
Moreover, we fully characterize how the parameter $\alpha$ depends on the underlying game characteristics. In particular, for the class of crowd motion games, we explicitly demonstrate how the resulting $\alpha$-Nash equilibria are shaped by the spatial distribution of the population, as well as the intensity and asymmetry of players’ interactions. This dependence is exemplified through the choice of the kernel function $K$ and the structure of the interaction weights $(q_{ij})_{i,j\in[N]}$ in the agents’ payoffs (1.2) (Theorem 6.1 and Corollary 6.1).
Finally, leveraging the algorithmic advantages of the $\alpha$-potential game paradigm as illustrated in [15], we develop an efficient policy-gradient algorithm (Algorithm 1) to minimize the $\alpha$-potential function and thereby construct an $\alpha$-Nash equilibrium. Through numerical experiments on crowd motion games, we showcase distinct emergent trajectories in both flocking and aversion dynamics.
@@
Let $T>0$. For each measurable function $\phi:[0,T]\to{\mathbb{R}}^{n}$, we define its $L^{2}$-norm $\|\phi\|_{L^{2}}=\left(\int_{0}^{T}|\phi(s)|^{2}{\mathrm{d}}s\right)^{1/2}$ with ||$\cdot$|| ⋅ | being the Euclidean norm, and for each $\phi:U\to{\mathbb{R}}^{m\times n}$ defined on a set $U$, we define its sup-norm $\|\phi\|_{L^{\infty}}=\sup_{u\in U}\|\phi(u)\|_{\textrm{sp}}$, with $\|\cdot\|_{\textrm{sp}}$ being the spectral norm of a matrix.
@@ 2.1. Mathematical Setup
Let $T>0$ be given terminal time, and $N,d,n,m,p\in\mathbb{N}$. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space which supports the following three mutually independent processes: a family of square integrable $d$-dimensional random variables $(\xi_{i})_{i=1}^{N}$, an $n$-dimensional Brownian motion $W=(W_{i})_{i=1}^{n}$, and a family of independent Poisson random measures $\eta=(\eta_{i})_{i=1}^{m}$ on $[0,T]\times\mathbb{R}^{p}_{0}$, where $\mathbb{R}^{p}_{0}\coloneqq{\mathbb{R}}^{p}\setminus\{0\}$ is equipped with its Borel $\sigma$-algebra. The random variables $(\xi_{i})_{i=1}^{N}$ represents the initial conditions of the system states, and the processes $W$ and $\tilde{\eta}$ represent the underlying system noises. We assume that each $\eta_{i}$ has a compensator $\nu_{i}({\mathrm{d}}z)\,{\mathrm{d}}t$, with $\nu_{i}$ being a $\sigma$-finite measure on $\mathbb{R}^{p}_{0}$ satisfying $\int_{\mathbb{R}^{p}_{0}}\min(1,|z|^{2})\,\nu_{i}({\mathrm{d}}z)<\infty,$ and define $\widetilde{\eta}_{i}({\mathrm{d}}t,{\mathrm{d}}z)=\eta_{i}({\mathrm{d}}t,{\mathrm{d}}z)-\nu_{i}({\mathrm{d}}z)\,{\mathrm{d}}t$ as the compensated Poisson random measure of $\eta_{i}$. Let $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ be the filtration generated by $(\xi_{i})_{i=1}^{N}$, $W$ and $\eta$, augmented with the $\mathbb{P}$-null sets.
We consider a stochastic differential game involving $N$ players, each employing open-loop control strategies defined as follows. For each $i\in[N]\coloneqq\{1,\ldots,N\}$, let $A_{i}\subset{\mathbb{R}}^{k}$ be player $i$’s action set, and let $\mathcal{A}_{i}$ be the set of player $i$’s admissible controls defined by
$\mathcal{A}_{i}\coloneqq\{u:\Omega\times[0,T]\to A_{i}\mid u\in\mathcal{H}^{2}({\mathbb{R}}^{k})\}.$ (2.1)
@@
${\mathrm{d}}X_{i,t}$ $=b_{i}(t)u_{i,t}{\mathrm{d}}t+\sigma_{i}(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\gamma_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad t\in(0,T];\quad X_{i,0}=\xi_{i},$ (2.2)
@@
$\inf_{u_{i}\in\mathcal{A}_{i}}J_{i}(\boldsymbol{u}),\quad\textnormal{with}\quad J_{i}(\boldsymbol{u})\coloneqq\mathbb{E}\left[\int_{0}^{T}f_{i}\left(t,\mathbf{X}_{t}^{\boldsymbol{u}},\boldsymbol{u}_{t}\right)\mathrm{d}t+g_{i}\left(\mathbf{X}_{T}^{\boldsymbol{u}}\right)\right],$ (2.3)
@@
Note the distinction between distributed games and distributed controls, the latter of which typically assumes that all players’ states are independent and that each player’s control depends only on their own state (see, e.g., [20]). In contrast, in distributed games players’ states and control processes can be correlated due to shared sources of randomness, such as correlated initial states and common components in the Brownian motions or the Poisson random measures
@@
${\mathrm{d}}\tilde{X}_{i,t}=\left(\tilde{a}_{i}(t)\,\tilde{X}_{i,t}+\tilde{b}_{i}(t)\,u_{i,t}\right){\mathrm{d}}t+\tilde{\sigma}_{i}(t)\,{\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\tilde{\gamma}_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad\tilde{X}_{i,0}^{u_{i}}=\xi_{i},$ (2.4)
@@
$X_{i,t}^{u_{i}}\coloneqq e^{-\int_{0}^{t}\tilde{a}_{i}(s)\,{\mathrm{d}}s}\,\tilde{X}_{i,t}^{u_{i}},\quad t\in[0,T],$
the state dynamics (2.4) can be transformed into the simpler form given in (2.2), with the state coefficients and cost functions adjusted by certain deterministic factors. A special case of (2.4) is the following controlled kinetic equation (see e.g, [33, 39]): for all $t\in[0,T]$,
[…]
@@
[…]
Under Assumption (H.1), for each ${\textbf{{u}}}\in\mathcal{H}^{2}({\mathbb{R}}^{kN})$, (2.2) admits a unique strong solution ${\textbf{X}}^{{\textbf{{u}}}}\in\mathcal{S}^{2}({\mathbb{R}}^{dN})$ (see [26, Theorem 3.1]), and (2.3) is well defined. For ease of exposition, we assume that the action set contains 0, but similar analyses can be extended to a non-empty convex action set (see e.g., [17, 18]).

### v2

@@ Abstract.
Motivated by game-theoretic models of crowd motion dynamics, this paper analyzes a broad class of distributed games with jump diffusions within the recently developed $\alpha$-potential game framework. We demonstrate that analyzing the $\alpha$-Nash equilibria reduces to solving a finite-dimensional control problem. Beyond the viscosity and verification characterizations for the general games, we examine explicitly and in detail how spatial population distributions and interaction rules influence the structure of $\alpha$-Nash equilibria in these distributed settings.
For crowd motion network games, we show that $\alpha=0$ for all symmetric interaction networks, and for asymmetric networks, we quantify the precise polynomial and logarithmic decays of $\alpha$ in terms of the number of players, the degree of the network, and the decay rate of interaction asymmetry. We also exploit the $\alpha$-potential game framework to analyze an $N$-player portfolio selection game under a mean–variance criterion. We show that this portfolio game constitutes a potential game and explicitly construct its Nash equilibrium. Our analysis allows for heterogeneous preference parameters, going beyond the mean-field interactions considered in the existing game literature.
Our theoretical results are supported by numerical implementations using policy gradient-based algorithms, demonstrating the computational advantages of the $\alpha$-potential game framework in computing Nash equilibria for general dynamic games.
@@
Potential game, Nash equilibrium, distributed game, policy gradient algorithm, game-theoretic motion planning, mean-variance portfolio game
@@ 1. Introduction
(1.1) ${\mathrm{d}}X_{i,t}$ $=b_{i}(t)u_{i,t}{\mathrm{d}}t+\sigma_{i}(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\gamma_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad t\in(0,T];\quad X_{i,0}=x_{i}\in{\mathbb{R}}^{d},$
@@
(1.2) $J_{i}({\textbf{{u}}})\coloneqq{\mathbb{E}}\left[\int_{0}^{T}\left(\ell_{i}(u_{i,t})+\frac{1}{N-1}\sum_{j=1,j\neq i}^{N}q_{ij}K(X^{u_{i}}_{i,t}-X^{u_{j}}_{j,t})\right)\,{\mathrm{d}}t+g_{i}(X^{u_{i}}_{i,T})\right],$
@@
The above crowd motion game is a special class of stochastic differential games which we name as distributed games; See Section 2 for the detailed formulation of these games. The term “distributed” refers to the characteristics of the game where each player’s dynamics evolve according to a controlled stochastic process that depends only on her own control, while her objective function may depend on the joint state and control profiles of all players (see also Remark 2.1). Such a framework has been used in a variety of applications where agents interact through their objectives but evolve independently in state, including distributed control of multi-agent systems and trajectory planning [27, 52, 53, 1], transportation and routing [4, 9, 29], as well as in energy markets and smart grids [35, 51, 44, 41, 55, 39]. The jump processes $(\tilde{\eta}_{j})_{j=1}^{m}$ in the state dynamics (1.1) capture sudden stochastic shocks and can be used to model neuronal spiking activity in neuroscience [38, 50], as well as stochastic demand shocks that introduce discontinuities in production and inventory management [33].
@@
Meanwhile, the recently introduced $\alpha$-potential game framework has shown significant promise for analyzing and solving general dynamic games, both from theoretical and algorithmic perspectives [16, 20, 17, 18, 26, 36]. The $\alpha$-potential game framework directly addresses finite-player games, unlike the conventional mean-field game approach, which relies on weak interactions among players or considers the limit as the number of players $N\to\infty$. Specifically, [16, 17] demonstrate that computing a Nash equilibrium can be reformulated as an optimization problem involving a single $\alpha$-potential function and the analysis of the parameter $\alpha$. Furthermore, [17] constructs a specific $\alpha$-potential function for general stochastic differential games and optimizes it by reformulating the problem as a conditional McKean–Vlasov control problem. As a result, [17] obtains an $\alpha$-Nash equilibrium via a Hamilton–Jacobi–Bellman (HJB) equation on the infinite-dimensional space of probability measures. However, beyond the LQ setting, analyzing this infinite-dimensional control problem and computing the $\alpha$-Nash equilibrium becomes highly challenging.
@@
In this paper, we develop the $\alpha$-potential game framework for distributed games with controlled jump-diffusion dynamics.
  • Within this framework, we show that the task of finding an $\alpha$-Nash equilibrium can be further reduced to solving a finite-dimensional control problem (Theorem 3.1), thereby avoiding the infinite-dimensional McKean–Vlasov control formulation given in [17]. A key step is to obtain a separation principle for the controlled state processes (Lemma 3.2). As a consequence, we obtain new Markovian characterizations of $\alpha$-Nash equilibria via a finite-dimensional HJB equation, utilizing both verification methods and viscosity solution theory (Theorems 4.1 and 4.3).
  • We precisely characterize how the parameter $\alpha$ depends on the underlying game structure, including the network governing player interactions, extending beyond the results in [17]. In particular, for games of the form (1.1)-(1.2), we explicitly demonstrate how the resulting $\alpha$-Nash equilibria are shaped by the choice of the kernel function $K$ and the intensity and asymmetry of the interaction weights $(q_{ij})_{i,j\in[N]}$ in the payoffs (Theorem 6.1 and Corollary 6.1). We prove that $\alpha=0$ for all symmetric interaction networks, and for asymmetric networks, we quantify the precise polynomial and logarithmic decays of $\alpha$ in terms of the number of players, the degree of the network, and the decay rate of interaction asymmetry.
  • Leveraging the finite-dimensional control formulation of the $\alpha$-potential function, we develop an efficient policy gradient algorithm (Algorithm 1) to compute an $\alpha$-Nash equilibrium. Through numerical experiments on crowd motion games, we showcase distinct emergent trajectories in both flocking and aversion dynamics.
  • We exploit the $\alpha$-potential game framework to analyze an $N$-player portfolio selection game under a mean–variance criterion. Players trade in a financial market with multiple risky assets, whose prices are driven by Brownian motions and Poisson random measures. Each player aims to maximize their total wealth relative to a weighted average of the other players’ wealth while penalizing variance. We show that this portfolio game constitutes a potential game and explicitly construct the corresponding Nash equilibrium (Theorem 7.1). Our analysis allows for heterogeneous preference parameters, going beyond the mean-field interactions considered in the existing game literature.
@@
Let $T>0$. For each measurable function $\phi:[0,T]\to{\mathbb{R}}^{n}$, we define its $L^{2}$-norm $\|\phi\|_{L^{2}}=\left(\int_{0}^{T}|\phi(s)|^{2}{\mathrm{d}}s\right)^{1/2}$ with $|\cdot|$ being the Euclidean norm, and for each $\phi:U\to{\mathbb{R}}^{m\times n}$ defined on a set $U$, we define its sup-norm $\|\phi\|_{L^{\infty}}=\sup_{u\in U}\|\phi(u)\|_{\textrm{sp}}$, with $\|\cdot\|_{\textrm{sp}}$ being the spectral norm of a matrix.
@@ 2.1. Mathematical Setup
Let $T>0$ be a given terminal time, and $N,d,n,m,p\in\mathbb{N}$. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space which supports the following three mutually independent processes: a family of square integrable $d$-dimensional random variables $(\xi_{i})_{i=1}^{N}$, an $n$-dimensional Brownian motion $W=(W_{i})_{i=1}^{n}$, and a family of independent Poisson random measures $\eta=(\eta_{i})_{i=1}^{m}$ on $0,T]\times\mathbb{R}^{p}_{0}$, where $\mathbb{R}^{p}_{0}\coloneqq{\mathbb{R}}^{p}\setminus\{0\}$ is equipped with its Borel $\sigma$-algebra. The random variables $(\xi_{i})_{i=1}^{N}$ represents the initial conditions of the system states, and the processes $W$ and $\tilde{\eta}$ represent the underlying system noises. We assume that each $\eta_{i}$ has a compensator $\nu_{i}({\mathrm{d}}z)\,{\mathrm{d}}t$, with $\nu_{i}$ being a $\sigma$-finite measure on $\mathbb{R}^{p}_{0}$ satisfying $\int_{\mathbb{R}^{p}_{0}}\min(1,|z|^{2})\,\nu_{i}({\mathrm{d}}z)<\infty,$ and define $\widetilde{\eta}_{i}({\mathrm{d}}t,{\mathrm{d}}z)=\eta_{i}({\mathrm{d}}t,{\mathrm{d}}z)-\nu_{i}({\mathrm{d}}z)\,{\mathrm{d}}t$ as the compensated Poisson random measure of $\eta_{i}$. Let $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ be the filtration generated by $(\xi_{i})_{i=1}^{N}$, $W$ and $\eta$, augmented with the $\mathbb{P}$-null sets.
We consider a stochastic differential game involving $N$ players, each employing open-loop control strategies defined as follows. [^1] Open-loop controls refer to strategies that are non-anticipative functions of the system noise and the initial states (see e.g., [7, Chapter 2]). We will prove in Section 4 that the resulting Nash equilibria admit a feedback representation. The $\alpha$-potential game framework can also be applied to games with closed-loop controls, where strategies are induced by policies that depend directly on the current system state [16, 11, 43]. In this setting, the characterization of $\alpha$ additionally depends on the regularity and structural properties of the closed-loop policies, and is therefore typically more involved than in the open-loop case. For each $i\in[N]\coloneqq\{1,\ldots,N\}$, let $A_{i}\subset{\mathbb{R}}^{k}$ be player $i$’s action set, and let $\mathcal{A}_{i}$ be the set of player $i$’s admissible controls defined by
(2.1) $\mathcal{A}_{i}\coloneqq\{u:\Omega\times[0,T]\to A_{i}\mid u\in\mathcal{H}^{2}({\mathbb{R}}^{k})\}.$
@@
(2.2) ${\mathrm{d}}X_{i,t}$ $=b_{i}(t)u_{i,t}{\mathrm{d}}t+\sigma_{i}(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\gamma_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad t\in(0,T];\quad X_{i,0}=\xi_{i},$
@@
(2.3) $\inf_{u_{i}\in\mathcal{A}_{i}}J_{i}(\boldsymbol{u}),\quad\textnormal{with}\quad J_{i}(\boldsymbol{u})\coloneqq\mathbb{E}\left[\int_{0}^{T}f_{i}\left(t,\mathbf{X}_{t}^{\boldsymbol{u}},\boldsymbol{u}_{t}\right)\mathrm{d}t+g_{i}\left(\mathbf{X}_{T}^{\boldsymbol{u}}\right)\right],$
@@
Note the distinction between distributed games and distributed controls, the latter of which typically assumes that all players’ states are independent and that each player’s control depends only on their own state (see, e.g., [24]). In contrast, in distributed games players’ states and control processes can be correlated due to shared sources of randomness, such as correlated initial states and common components in the Brownian motions or the Poisson random measures.
@@
(2.4) ${\mathrm{d}}\tilde{X}_{i,t}=\left(\tilde{a}_{i}(t)\,\tilde{X}_{i,t}+\tilde{b}_{i}(t)\,u_{i,t}\right){\mathrm{d}}t+\tilde{\sigma}_{i}(t)\,{\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\tilde{\gamma}_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad\tilde{X}_{i,0}^{u_{i}}=\xi_{i},$
@@
$X_{i,t}^{u_{i}}\coloneqq A_{i}(t)\tilde{X}_{i,t}^{u_{i}},\quad t\in[0,T],$
with $\frac{{\mathrm{d}}}{{\mathrm{d}}t}A_{i}(t)=-A_{i}(t)\tilde{a}_{i}(t)$ and $A_{i}(0)=I_{d}$, the state dynamics (2.4) can be transformed into the simpler form given in (2.2), with the state coefficients and cost functions adjusted by certain deterministic factors. A special case of (2.4) is the following controlled kinetic equation (see e.g, [40, 47]): for all $t\in[0,T]$,
$\left\{\begin{aligned} {\mathrm{d}}x_{i,t}&=v_{i,t}{\mathrm{d}}t,\quad x_{i,0}=x_{i},\\ {\mathrm{d}}v_{i,t}&=u_{i,t}{\mathrm{d}}t+\tilde{\sigma}_{i}(t)\,{\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\tilde{\gamma}_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad v_{i,0}=v_{i},\end{aligned}\right$.
@@
    For all $t\in[0,T]$, ${\mathbb{R}}^{dN}\times{\mathbb{R}}^{kN}\ni(x,a)\mapsto(f_{i}(t,x,a),g_{i}(x))\in{\mathbb{R}}\times{\mathbb{R}}$ is twice continuously differentiable, $0,T]\ni t\mapsto\big(f_{i}(t,0,0),\partial_{(x,a)}f_{i}(t,0,0)\big)\in{\mathbb{R}}\times{\mathbb{R}}^{(d+k)N}$ is bounded, and the second-order derivatives $\partial^{2}_{xx}f_{i}$, $\partial^{2}_{xa}f_{i}$, $\partial^{2}_{aa}f_{i}$, and $\partial^{2}_{xx}g_{i}$ are bounded (uniformly in $(t,x,a)$).
Under Assumption (H.1), for each ${\textbf{{u}}}\in\mathcal{H}^{2}({\mathbb{R}}^{kN})$, (2.2) admits a unique strong solution ${\textbf{X}}^{{\textbf{{u}}}}\in\mathcal{S}^{2}({\mathbb{R}}^{dN})$ (see [30, Theorem 3.1]), and (2.3) is well defined. For ease of exposition, we assume that the action set contains 0, but similar analyses can be extended to a non-empty convex action set (see e.g., [17, 20]).